OpenStudy (rayford):

what is the simplified form of x^2-25/x+5

OpenStudy (ash2326):

@rayford we are given \[\frac{x^2-25}{x+5}\] This could be written as \[\frac{x^2-5^2}{x+5}\] We could use here \[a^2-b^2=(a+b)(a-b) \] Can you try from here?

OpenStudy (ash2326):

@rayford do you understand till here?

OpenStudy (rayford):

i'm sorry i was called for something. i'm sorta understanding until here..may you continue?

OpenStudy (ash2326):

Read my post and try to use the relation \[a^2-b^2=(a+b)(a-b)\] to simplify \(x^2-5^2\) Once you do that, you'd be able to simplify easily

OpenStudy (rayford):

x + 5, with the restriction x -5

OpenStudy (ash2326):

@rayford could you check again, there is a small mistake?

OpenStudy (rayford):

hmmm....the only conclusion i could come down to is probably x + 5, with the restriction x 5

OpenStudy (ash2326):

\[\frac{x^2-5^2}{x+5}\ \ =\ \ \frac{\cancel{(x+5)}(x-5)}{\cancel{x+5}}\] \[x-5\]

OpenStudy (rayford):

whats the restriction though?

OpenStudy (rayford):

5 or -5?

OpenStudy (ash2326):

it's on \[x+5\] It can't be zero. so \(x\ne -5\)

OpenStudy (rayford):

life saver man...thank you.

OpenStudy (ash2326):

Do you understand this?